undeadknight26 (undeadknight26):

Medal! f(x)=a(b)x, where a is the y-intercept and b is the base of the exponential expression.

hartnn (hartnn):

you mean \(\large f(x)=ab^x\) ? what you need to do with this ?

undeadknight26 (undeadknight26):

a question for school....

hartnn (hartnn):

i mean i don't see a question anywhere .....? what you need to do with f(x) ?

undeadknight26 (undeadknight26):

I have no idea thats all they gave me...

hartnn (hartnn):

can you post a screenshot of the page ?

undeadknight26 (undeadknight26):

this is what they said. Exponential functions • f(x)=a(b)x, where a is the y-intercept and b is the base of the exponential expression. • f(x)=P(1+r)x, where P is the principal amount and r is the rate of change in decimal form.

hartnn (hartnn):

yes, so they have given example of exponential equations. what they want you to do with it ?

hartnn (hartnn):

see, if there is no question, there will be no answer too

undeadknight26 (undeadknight26):

Lol true.

hartnn (hartnn):

so thats just a review :) read it and learn. no questions, no answers :)

hartnn (hartnn):

thanks for sharing knowledge :)

undeadknight26 (undeadknight26):

just noticed that makes me look stupid...*facepalm*

undeadknight26 (undeadknight26):

Thanks for bearing with me man!

hartnn (hartnn):

you're welcome ^_^